# Neural network learns low-dimensional polynomials with SGD near the information-theoretic limit

### v1

@@ 4.2 SGD with Batch Reuse Implements Polynomial Transformation
$\eta\bigg{[}y\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle+\sum_{i=0}^{C_{\sigma}-1}\underbrace{(\eta\|\boldsymbol{x}\|^{2})^{i}y^{i+1}(i!)^{-1}(\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle))^{i}\sigma^{(i+1)}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle}_{=:\psi_{i}}\bigg{]}.$ (4.2)
We take $\eta\leq c_{\eta}d^{-1}$ with a small constant $c_{\eta}$ so that $\eta\|\boldsymbol{x}\|^{2}\ll 1$. Crucially, the strength of each term in (4.2) can vary depending on properties of the link function $\sigma_{*}$, which is unknown. Hence a careful analysis is required to ensure that the suitable monomial transformation is always singled out from the gradient update. We therefore divide our analysis into four cases (for simplicity we present the noiseless setting below).
  • (I)
    If $\mathrm{IE}(\sigma_{*})=1$. All terms in the summation in (4.2) with $i\geq 2$ decay as fast as $\eta\|\boldsymbol{x}\|^{2}\ll 1$. On the other hand, the expectation of $y\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle$ is roughly $\alpha_{1}\beta_{1}\gtrsim 1$. Therefore we may isolate the informative term $y\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle$. This case is discussed in Section B.3.3.
  • (II)
    Else if $\mathrm{IE}((\sigma_{*})^{I})=1$ for some $2\leq I\leq C_{\sigma}$. Let $I$ be the lowest degree of monomial transformation such that $\mathrm{IE}((\sigma_{*})^{I})=1$. Since $\sigma_{*},\cdots,\sigma_{*}^{I-1}$ have information exponent larger than $1$, expectations of $y\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle$ and $\psi_{i}\ (i=2,\cdots,I-2)$ scales as $\langle\boldsymbol{\theta},\boldsymbol{\omega}\rangle\asymp d^{-1/2}$. For $i=I-1$, because $\sigma_{*}^{I}$ has information exponent $1$,
    $\mathbb{E}[\psi_{I-1}]=\,$ $\mathbb{E}\big{[}(\eta\|\boldsymbol{x}\|^{2})^{I-1}y^{I}((I-1)!)^{-1}(\sigma^{(1)}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle))^{I-1}\sigma^{(I)}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle\big{]}$ (4.3)
@@
    For $i\geq I$, $\psi_{i}$ decays as $c_{\eta}^{I}$, which is smaller than the scale of $\psi_{I-1}\asymp c_{\eta}^{I-1}$. Hence the term $\psi_{I-1}\gtrsim c_{\eta}^{I-1}$ is singled out. This case is discussed in Section B.3.2.
  • (III)
    Else if $\mathrm{IE}(\sigma_{*})=2$. We have $\mathbb{E}[y\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle]\approx 2\alpha_{2}\beta_{2}\langle\boldsymbol{\theta},\boldsymbol{\omega}\rangle$. Also, for $i\geq 2$, since $\sigma_{*}^{2},\cdots,\sigma_{*}^{C_{\sigma}}$ have information exponent at least $2$, expectation of $\psi_{i}$ is roughly of order $(\eta\|\boldsymbol{x}\|^{2})^{i}\langle\boldsymbol{\theta},\boldsymbol{\omega}\rangle$. Therefore, the term $y\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle$ is singled out, and the expectation scales as $\alpha_{2}\beta_{2}d^{-1/2}$ at initialization. This case is discussed in Section B.3.3.
  • (IV)
    Else $\mathrm{IE}((\sigma_{*})^{I})=2$ for some $2\leq I\leq C_{\sigma}$. In this case, since $\sigma_{*},\cdots,\sigma_{*}^{I-1}$ have information exponent larger than $2$, expectations of $y\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle$ and $\psi_{i}\ (i=2,\cdots,I-2)$ are at most $(\langle\boldsymbol{\theta},\boldsymbol{\omega}\rangle)^{2}\asymp d^{-1}$. And at $i=I-1$, because $\sigma_{*}^{I}$ has information exponent $2$,
    $\mathbb{E}[\psi_{I-1}]=\,$ $\mathbb{E}\big{[}(\eta\|\boldsymbol{x}\|^{2})^{I-1}y^{I}((I-1)!)^{-1}(\sigma^{(1)}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle))^{I-1}\sigma^{(I)}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle\big{]}$ (4.5)
[…]
    As for $i\geq I$, because $\sigma_{*}^{i}$ has information exponent larger than $1$ for $I+1\leq i\leq C_{\sigma}$, $\psi_{i}$ decays as $c_{\eta}^{I}\langle\boldsymbol{\theta},\boldsymbol{\omega}\rangle$, which is smaller than $\psi_{I-1}\asymp c_{\eta}^{I-1}\langle\boldsymbol{\theta},\boldsymbol{\omega}\rangle$. Thus, the term $\psi_{I-1}$ is dominating, whose scale is roughly $c_{\eta}^{I-1}d^{-1/2}$ at initialization. This case is discussed in Section B.3.1.
Why interpolation is required.
[…]
Combining the four cases yields the following lemma on the evolution of alignment.
@@
Under the assumptions per Theorem 1, one of the following holds:
  • (i)
@@
    $\langle\boldsymbol{\theta},\boldsymbol{w}_{j}^{2(t+1)}\rangle\geq\langle\boldsymbol{\theta},\boldsymbol{w}_{j}^{2t}\rangle+\eta^{2t}_{j}(1-\xi_{j}^{2(t+1)})\gamma+\eta^{2t}_{j}(1-\xi_{j}^{2(t+1)})\nu^{2t}_{j}.$
  • (ii)
    $\langle\boldsymbol{\theta},\boldsymbol{w}_{j}^{2(t+1)}\rangle\geq\langle\boldsymbol{\theta},\boldsymbol{w}_{j}^{2t}\rangle+\eta^{2t}_{j}(1-\xi_{j}^{2(t+1)})\gamma\langle\boldsymbol{\theta},\boldsymbol{w}_{j}^{2t}\rangle+\eta^{2t}_{j}(1-\xi_{j}^{2(t+1)})\nu^{2t}_{j}.$
Here $\gamma\gtrsim c_{\eta}^{I-1}$ is a constant that depends on $\sigma_{*}$ and $\nu^{2t}_{j}$ is a mean-zero noise.
For (i), taking expectation immediately yields that weak recovery is achieved within $(\eta(1-\xi)\gamma)^{-1}=\tilde{O}(d)$ steps. For (ii), $\langle\boldsymbol{\theta},\boldsymbol{w}_{j}^{2t}\rangle=:\kappa^{t}$ can be approximated by a differential equation $\frac{\mathrm{d}\kappa^{t}}{\mathrm{d}t}=\eta(1-\xi)\gamma\kappa^{t}$. Solving this yields $\kappa^{t}=\kappa^{0}\exp(\eta(1-\xi)\gamma t)\approx d^{-\frac{1}{2}}\exp(\eta(1-\xi)\gamma t)$, and weak recovery is obtained within $t\lesssim(\eta(1-\xi)\gamma)^{-1}\cdot\log d=\tilde{O}(d)$ steps, similar to the analysis in [BAGJ21].
4.3 Analysis of Phase II and Statistical Guarantees
Once strong recovery is achieved for the first-layer parameters, we turn to Phase II and optimize the second-layer with $\ell_{2}$ regularization. Since the objective is strongly convex, gradient-based optimization can efficiently minimize the empirical loss. The learnability guarantee follows from standard analysis analogous to that in [AAM22, DLS22, BES+22], where we construct a “certificate” second-layer $\boldsymbol{a}^{*}\in\mathbb{R}^{N}$ that achieves small loss and small norm:
from which the population loss of the regularized empirical risk minimizer can be bounded via standard Rademacher complexity argument. To construct such a certificate, we make use of the random bias units $\{b_{j}\}_{j=1}^{N}$ to approximate the link function $\sigma_{*}$ as done in [DLS22, BBSS22, OSSW24].
@@ 5 Conclusion and Future Directions
In this work we showed that a two-layer neural network (3.1) trained by SGD with reused batch can learn arbitrary single-index polynomials up to $\varepsilon$ population error using $n=\tilde{O}(d\varepsilon^{-2})$ samples and compute. Our analysis is based on the observation that by reusing the same minibatch twice in the gradient computation, a non-correlational term arises in the SGD update that transforms the labels (despite the loss function is not modified). Specifically, following the definition in [DPVLB24], we know that polynomial $\sigma_{*}$ has generative exponent at most $2$, which implies the existence of nonlinear transformation $\mathcal{T}:\mathbb{R}\to\mathbb{R}$ such that the information exponent $p_{*}$ becomes at most $2$, i.e.,
$\mathbb{E}[\mathcal{T}(y)\mathsf{He}_{i}(\langle\boldsymbol{x},\boldsymbol{\theta}\rangle)]\neq 0,\quad\text{for $i=1$ or $2$.}$
We show that restricting $\mathcal{T}$ to be polynomial is sufficient, and such transformation can be extracted by Taylor-expanding the SGD update. Then we show via careful analysis of the trajectory that strong recovery and low population error can be achieved under suitable activation function.
Future directions.
First, our analysis only handles link functions with generative exponent $p_{*}\leq 2$; while this covers arbitrary polynomial $\sigma_{*}$ analogous to [CM20], it is interesting to examine whether SGD with reused batch can learn targets with $p_{*}\geq 3$ with a sample complexity matching the SQ lower bound. It is also possible that ERM algorithms on i.i.d. data can achieve a statistical complexity beyond the SQ lower bound due to non-adversarial noise [DH21, DH24]; such mechanism is not exploited in our current analysis. Additional interesting directions include extension to multi-index [BAGJ22, BBPV23, CWPPS23, Gla23], hierarchical polynomials [NDL23], and additive models [OSSW24]. Lastly, the SGD algorithm that we employ requires a layer-wise training procedure and a specific batch reuse schedule; one may therefore ask if standard multi-pass SGD training of all parameters simultaneously (as reported in Figure 1) also achieves the same statistical efficiency.
@@ Acknowledgements
The authors thank Gerard Ben Arous, Joan Bruna, Alex Damian, Marco Mondelli, and Eshaan Nichani for the discussions and feedback on the manuscript. JDL acknowledges support of the ARO under MURI Award W911NF-11-1-0304, NSF CCF 2002272, NSF IIS 2107304, NSF CIF 2212262, ONR Young Investigator Award, and NSF CAREER Award 2144994. KO was partially supported by JST ACT-X (JPMJAX23C4). TS was partially supported by JSPS KAKENHI (24K02905) and JST CREST (JPMJCR2015).
@@ A.1 Proof for Even Functions $(i)$
$\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\bigg{(}\frac{f(t)}{2\tau(f)}\bigg{)}^{i}(t^{2}-1)\bigg{]}.$ (A.1)
@@
$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[\tau(f)\leq|f(t)|<2\tau(f)]\bigg{(}\frac{f(t)}{2\tau(f)}\bigg{)}^{i}(t^{2}-1)\bigg{]}$ (A.3)
$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[|f(t)|<\tau(f)]\bigg{(}\frac{f(t)}{2\tau(f)}\bigg{)}^{i}(t^{2}-1)\bigg{]}$ (A.4)
$\geq$ $~{}\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[|f(t)|\geq 2\tau(f)]\bigg{(}\frac{2\tau(f)}{2\tau(f)}\bigg{)}^{i}(2^{2}-1)\bigg{]}$ (A.5)
$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[\tau(f)\leq|f(t)|<2\tau(f)]\bigg{(}\frac{f(t)}{2\tau(f)}\bigg{)}^{i}(2^{2}-1)\bigg{]}$ (A.6)
$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[|f(t)|<\tau(f)]\bigg{(}\frac{\tau(f)}{2\tau(f)}\bigg{)}^{i}(0^{2}-1)\bigg{]}$ (A.7)
$\geq$ $~{}3\mathbb{P}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[|f(t)|\geq 2\tau(f)]-2^{-i}.$ (A.8)
@@
$2H(f^{i};2)^{2}\leq\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}(f(t))^{2i}\big{]}\leq(2i-1)^{iq}\big{(}\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}f(t)^{2}\big{]}\big{)}^{i}=(2i-1)^{iq}.$ (A.10)

### v2

@@ 4.2 SGD with Batch Reuse Implements Polynomial Transformation
$\eta\bigg{[}\textstyle y\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle+\sum_{i=0}^{C_{\sigma}-1}\underbrace{\textstyle(\eta\|\boldsymbol{x}\|^{2})^{i}y^{i+1}(i!)^{-1}(\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle))^{i}\sigma^{(i+1)}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle}_{=:\psi_{i}}\bigg{]}.$ (4.2)
We take $\eta\leq c_{\eta}d^{-1}$ with a small constant $c_{\eta}$ so that $\eta\|\boldsymbol{x}\|^{2}\ll 1$ with high probability. Crucially, the strength of each term in (4.2) can vary depending on properties of the unknown link function $\sigma_{*}$. Hence a careful analysis is required to ensure that a suitable monomial transformation is always singled out from the gradient. We establish the following lemma on the evolution of alignment.
Lemma 7.
Under the assumptions per Theorem 1, the following holds for generative exponent $p_{*}=1,2$:
$\langle\boldsymbol{\theta},\boldsymbol{w}^{2(t+1)}\rangle\geq\langle\boldsymbol{\theta},\boldsymbol{w}^{2t}\rangle+c_{\eta}^{I}c_{\xi}c_{\sigma}d^{-\frac{p_{*}}{2}\lor 1}(\kappa^{2t})^{p_{*}-1}+c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}\lor 1}\nu^{2t}.$ (4.3)
See Lemma 16 for the formal version. For $p_{*}=1$, taking expectation immediately yields that weak recovery within $(\eta(1-\xi)\gamma)^{-1}=O(d)$ steps. For $p_{*}=2$, $\langle\boldsymbol{\theta},\boldsymbol{w}_{j}^{2t}\rangle=:\kappa^{t}$ can be approximated by a differential equation $\frac{\mathrm{d}\kappa^{t}}{\mathrm{d}t}=\eta(1-\xi)\gamma\kappa^{t}$. Solving this yields $\kappa^{t}=\kappa^{0}\exp(\eta(1-\xi)\gamma t)\approx d^{-\frac{1}{2}}\exp(\eta(1-\xi)\gamma t)$, and weak recovery is obtained within $t\lesssim(\eta(1-\xi)\gamma)^{-1}\cdot\log d=O(d\log d)$ steps, similar to the analysis in [BAGJ21].
Why interpolation is needed.
@@
4.3 Analysis of Phase II and Statistical Guarantees
Once strong recovery is achieved for the first-layer parameters, we turn to Phase II and optimize the second-layer with $\ell_{2}$ regularization. Since the objective is strongly convex, gradient-based optimization can efficiently minimize the empirical loss. In Appendix B.6, the learnability guarantee follows from standard analysis analogous to that in [AAM22, DLS22, BES+22], where we construct a “certificate” second-layer $\boldsymbol{a}^{*}\in\mathbb{R}^{N}$ that achieves small loss and small norm:
from which the population loss of the regularized empirical risk minimizer can be bounded via standard Rademacher complexity argument. To construct such a certificate, we make use of the random bias units $\{b_{j}\}_{j=1}^{N}$ to approximate the link function $\sigma_{*}$ as done in [DLS22, BBSS22, OSSW24a].
5 Beyond Polynomial Link Functions
Thus far we have shown that for polynomial single-index target functions (which satisfy $p_{*}\leq 2$), SGD with data reuse can implement a polynomial transformation to the labels that reduces the information exponent to at most 2; consequently, the trained two-layer neural network can achieve small generalization error with $n=d\operatorname{polylog}(d)$ samples. However, as shown in [DPVLB24], there exists (non-polynomial) $\sigma_{*}$ with generative exponent $p_{*}>2$ (i.e., label transformations cannot lower the information exponent to $2$) and thus not learnable by SQ algorithms in linear sample complexity.
Nevertheless, for a single-index model with generative exponent $p_{*}$, we know there exists an “optimal” label transformation that reduces the information exponent to $p_{*}$. If SGD can make use of such transformation, then from the arguments in [BAGJ21], it is natural to conjecture that a sample size of $n\simeq d^{p_{*}-1}$ is sufficient. In this section we confirm this intuition by proving that SGD with data reuse (Algorithm 1) indeed matches this complexity. The following lemma is an analogue of Proposition 6 stating that polynomial transformations are sufficient to lower the information exponent.
@@ 5 Beyond Polynomial Link Functions
Given link function $\sigma_{*}$ with generative exponent $p_{*}\in\mathbb{N}_{+}$. Suppose we can take an orthonormal polynomial basis $\{\phi_{k}\}_{k}$ for the space $L^{2}(P_{y})$ with inner product $\langle f,g\rangle=\mathbb{E}_{y=\sigma_{*}(z)}[f(y)g(y)]$. Then there exists some degree of transformation $I\in\mathbb{N}_{*}$ such that $\mathrm{IE}(\sigma_{*}^{I})=p_{*}$.
We outline the differences and additional technical challenges to handle the $\mathrm{GE}(\sigma_{*})>2$ setting.
  • For general $L^{2}$ link functions $\sigma_{*}$, we can no longer make use of the compactness argument (see proof of Proposition 6) to upper bound the degree of monomial transformation. Hence in Lemma 8 we do not state a uniform upper bound on the required degree $I$, unlike the polynomial setting.
  • Any link function with $p_{*}>2$ cannot be polynomial, and hence we cannot achieve low generalization error using a neural network with polynomial nonlinearity. We therefore need to use an activation function with universal function approximation ability.
5.1 Sample Complexity for Weak Recovery
We first show that Algorithm 1 achieves weak recovery (i.e., nontrivial overlap with the ground truth $\boldsymbol{\theta}$) with a complexity governed by the generative exponent of the link function $p_{*}=\mathrm{GE}(\sigma_{*})$. Similar to Section 3.2, we make use of randomized activation functions to ensure the desired label transformation is encoded — we defer the conditions on the student activation to Appendix B.1.2. Similar to Theorem 1, we focus on the subset of neurons with large initial overlap, and activation satisfying the assumptions in Appendix B.1.2 (these conditions are met by $\tilde{\Omega}(1)$ fraction of neurons).
Proposition 9.
Suppose the link function $\sigma_{*}$ has generative exponent $p_{*}$, and let $I\in\mathbb{N}_{+}$ be the smallest degree of monomial transformation that lowers the information exponent to $p_{*}$ (i.e., $\mathrm{IE}(\sigma_{*}^{I})=p$). We can find a student activation function $\sigma$ depending only on $p,p_{*}$ and $I$, such that if we take $\eta^{2t},\eta^{2t+1}=c_{\eta}Nd^{-1}$, $\xi^{2(t+1)}=1-c_{\xi}d^{-\nicefrac{{(p_{*}-2)_{+}}}{{2}}}$ for small $c_{\eta},c_{\xi}=o_{d}(1)$, and set
$T_{1,1}\simeq c_{\xi}^{-1}\begin{cases}d&(\text{if $p_{*}=1$})\\ d(\log d)&(\text{if $p_{*}=2$})\\ d^{p_{*}-1}&(\text{if $p_{*}\geq 3$}),\end{cases}$ (5.1)
then if the initial alignment $\langle\boldsymbol{w}^{0},\boldsymbol{\theta}\rangle\geq 2c_{\eta}^{-1}d^{-\nicefrac{{1}}{{2}}}$, there exists $\tau_{*}\leq T_{1,1}$ such that for all $\tau\geq\tau_{*}$,
@@ 5.1 Sample Complexity for Weak Recovery
$\langle\boldsymbol{w}^{2\tau},\boldsymbol{\theta}\rangle\geq\tilde{\Theta}(1),\quad\text{with probability $1-o_{d}(1)$.}$ (5.2)
Proposition 9 is a generalization of Theorem 1 beyond polynomial $\sigma_{*}$ (the proof of both results are presented in Appendix B.3,B.4), and can be interpreted as an SQ counterpart to [BAGJ21]: we establish a sufficient sample size of $n\simeq d^{(p_{*}-1)\vee 1}$ for Algorithm 1 to exit the search phase, which is parallel to the $n\simeq d^{(p-1)\vee 1}$ rate for one-pass SGD (note that our rates are slightly sharper due to logarithmic factors removed, since $c_{\xi}^{-1}$ can grow arbitrarily slowly with $d$). For high generative exponent $\sigma_{*}$ with $p_{*}>2$, we no longer match the information theoretically optimal sample complexity $n\asymp d$, which is consistent with the computational-to-statistical gap observed in [DPVLB24].
5.2 Generalization Error Guarantee
After Phase I of Algorithm 1, we learn the unknown link function $\sigma_{*}$ via training the second-layer. To approximate non-polynomial functions, we introduce a ReLU component in the student nonlinearity $\sigma$ (see Lemma 12 for discussions), and make use of the approximation result for the (univariate) ReLU kernel in [BBSS22], which handles general $\sigma_{*}$ whose second derivative has bounded 4th moment. Combining the above, we arrive at the following end-to-end guarantee for learning single-index models with arbitrary generative exponent using SGD training of neural network.
Proposition 10 (Informal).
Suppose the link function $\sigma_{*}$ has generative exponent $p_{*}\in\mathbb{N}_{*}$ and satisfies $\sigma_{*},\sigma_{*}^{\prime\prime}\in L^{4}(\gamma)$. For appropriately chosen activation function $\sigma$ (see Appendix B.1.2), a neural network (3.1) with $N=\tilde{\Theta}(1)$ neurons optimized by Algorithm 1 achieves small population loss $\mathbb{E}_{\boldsymbol{x}}[(f_{\hat{\boldsymbol{\Theta}}}(\boldsymbol{x})-f_{*}(\boldsymbol{x}))^{2}]=o_{d,\mathbb{P}}(1)$ with a sample complexity of $n=\tilde{\Theta}(d^{(p_{*}-1)\vee 1})$.
See Appendix B.6 for the full statement with $\varepsilon$ dependence. This proposition confirms that the sample complexity for weak recovery (Proposition 9) is the bottleneck in single-index learning, as the total sample size required for Algorithm 1 to achieve low test error also scales with $d^{(p_{*}-1)\vee 1}$.
@@ 6 Conclusion and Future Directions
We showed that a two-layer neural network (3.1) trained by SGD with reused batch can learn single-index model (with generative exponent $p_{*}$) using $n\simeq d^{(p_{*}-1)\vee 1}$ samples and compute; in particular, when the link function $\sigma_{*}$ is polynomial, we established a sample complexity of $n=\tilde{O}(d\varepsilon^{-2})$ to achieve $\varepsilon$ population loss, which is almost information theoretically optimal. Our analysis is based on the observation that by reusing the same training data twice in the gradient computation, a non-correlational term arises in the SGD update that transforms the labels (despite the loss function not modified). We proved that monomial transformations that lower the information exponent of $\sigma_{*}$ can be extracted by Taylor-expanding the SGD update; then we showed via careful analysis of the trajectory that strong recovery and low population loss is achieved under suitable activation function.
Interesting future directions include extension to multi-index models [BAGJ22, BBPV23, CWPPS23], hierarchical target functions [AZL19, NDL23], and in-context learning [OSSW24b]. Also, the SGD algorithm that we employ requires a layer-wise training procedure and a specific batch reuse schedule; one may therefore ask if standard multi-pass SGD training of all parameters simultaneously [Gla23] (as reported in Figure 1) also achieves the same statistical efficiency.
@@ Acknowledgements
The authors thank Gerard Ben Arous, Joan Bruna, Alex Damian, Marco Mondelli, and Eshaan Nichani for the discussions and feedback on the manuscript. JDL acknowledges support of the ARO under MURI Award W911NF-11-1-0304, NSF CCF 2002272, NSF IIS 2107304, NSF CIF 2212262, ONR Young Investigator Award, and NSF CAREER Award 2144994. KO was partially supported by JST ACT-X (JPMJAX23C4). TS was partially supported by JSPS KAKENHI (24K02905) and JST CREST (JPMJCR2015). This research is unrelated to DW’s work at xAI.
@@ A.1 Proof for Even Functions $(i)$
$\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\bigg{(}\frac{f(t)}{2\tau(f)}\bigg{)}^{i}(t^{2}-1)\bigg{]}.$ (A.1)
@@
$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[\tau(f)\leq|f(t)|<2\tau(f)]\bigg{(}\frac{f(t)}{2\tau(f)}\bigg{)}^{i}(t^{2}-1)\bigg{]}$ (A.3)
$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[|f(t)|<\tau(f)]\bigg{(}\frac{f(t)}{2\tau(f)}\bigg{)}^{i}(t^{2}-1)\bigg{]}$ (A.4)
$\geq$ $~{}\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[|f(t)|\geq 2\tau(f)]\bigg{(}\frac{2\tau(f)}{2\tau(f)}\bigg{)}^{i}(2^{2}-1)\bigg{]}$ (A.5)
$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[\tau(f)\leq|f(t)|<2\tau(f)]\bigg{(}\frac{f(t)}{2\tau(f)}\bigg{)}^{i}(2^{2}-1)\bigg{]}$ (A.6)
$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[|f(t)|<\tau(f)]\bigg{(}\frac{\tau(f)}{2\tau(f)}\bigg{)}^{i}(0^{2}-1)\bigg{]}$ (A.7)
$\geq$ $~{}3\mathbb{P}_{t\sim\mathcal{N}(0,1)}[|f(t)|\geq 2\tau(f)]-2^{-i}.$ (A.8)
@@
$2H(f^{i};2)^{2}\leq\mathbb{E}_{t\sim\mathcal{N}(0,1)}\big{[}(f(t))^{2i}\big{]}\leq(2i-1)^{iq}\big{(}\mathbb{E}_{t\sim\mathcal{N}(0,1)}\big{[}f(t)^{2}\big{]}\big{)}^{i}=(2i-1)^{iq}.$ (A.10)